# The Nontrivial Vacuum Structure of an Extended $t\overline{t}$ BEH (Higgs) Bound state

### v1

@@ V Summary
We have proposed a nontrivial vacuum for the natural top condensation theory. The vacuum is manifestly Lorentz invariant, composed of a collective Lorentz invariant sum over internal wave-functions, $\Phi=(1/\sqrt{N})\sum_{\omega}\phi_{\omega}(r^{\mu})$. Though each internal wave-function, $\phi_{\omega}$, is independent of relative time, $\tau$, they each have dependence upon the arrow of relative time, $\omega_{\mu}$, and are solutions to the Schrödinger-Klein-Gordon equation with eigenvalue $|\mu^{2}|$. The sum over $\phi_{\omega}$ makes the collective field, $\Phi(r^{\mu})$, Lorenz invariant and completely independent of $\omega_{\mu}$.
@@
$H^{\prime}(X)\Phi(r)=H(X)\widetilde{\Phi}(r)\rightarrow\exp(i\pi^{a}(X)\tau^{a}/2v_{weak})\biggl(\begin{array}[]{c}v_{weak}\widetilde{\Phi}(r^{\mu})+h(X,r^{\mu})\\ 0\end{array}\biggr)$ (68)
where $v_{weak}=|\mu|/\sqrt{2}$ and the observed BEH boson is:
$h(X,r^{\mu})=h(X)\widetilde{\Phi}(r^{\mu})$ (69)
@@
We comment briefly on some phenomenological implications of this theory. If we go beyond the pointlike limit in, e.g., the BEH-Yukawa interaction, we will have Lorentz invariant $O(1/M_{0}^{2})$ corrections. Before taking the pointlike limit of the Yukawa interaction, as in eq.77), we can rewrite this in the broken phase as:
$\!\!\!\!\!\!\!\!\!\!S_{Y}^{\prime}=m_{top}\int\!\!d^{4}Xd^{4}r\;[\overline{\psi}_{L}(X\!+\!r)\psi_{R}(X\!-\!r)]_{f}(M_{0}^{2}D_{F}(2r^{\mu}))\;F(r^{\mu})+h.c.\qquad F=\frac{\widetilde{\Phi}(r)}{\widetilde{\Phi}(0)}\sim 1+a_{1}r^{2}+$... (70)
An expansion in $r^{\mu}$ in all terms in the integrand generates a tower of assorted higher dimension operators such as:
$\frac{m_{t}}{M_{0}^{2}}\biggl(1+\frac{h(X)}{\sqrt{2}v}\biggr)\biggl([\overline{b}_{L}(X)D^{2}t_{R}(X)]+[\overline{t}_{L}(X)D^{2}t_{R}(X)]+...\biggr)+h.c.$ (71)
$\overline{t}\rightarrow b+W+(g,\gamma,Z))\qquad\text{and,}\qquad\overline{t}\rightarrow t+(g,\gamma,Z))$ (72)
$(\ell^{+}\ell^{-})\rightarrow(\gamma^{*},Z^{*})\rightarrow t+b+(W,g,\gamma,Z))$ (73)
@@
Note that bilocality of the wave-function is an important naturalness constraint. One might be tempted to, e.g., loop the Yukawa interaction and argue for a problematic large correction to the BEH mass $\propto-\overline{g}_{Y}^{2}N_{c}M_{0}^{2}$. This would lead to the false conclusion that the effective theory is “unnatural.” The loop actually generates an enhancing correction to the bilocal potential, i.e., $\sim g_{Y}^{2}\delta^{4}(2r^{\mu})/M_{0}^{2}\sim g_{Y}^{2}D_{F}(2r^{\mu})$ for large $M_{0}$ chris2 , and we show that this leads to a “critical amplification” of the effective 4-fermion coupling, $\overline{g}^{2}_{0}$, where: $\overline{g}_{0}^{2}=g_{0}^{2}\left(1-{g_{0}^{2}N_{c}}/{8\pi^{2}}\right)^{-1}.$ Hence, while $\overline{g}_{0}^{2}$ is supercritical, the underlying topcolor coupling, $g_{0}^{2}$, is smaller and subcritical.
@@
In conclusion, we emphasize this theory is natural and manifestly Lorentz invariant. The approximate scale symmetry near critical coupling provides the custodial symmetry of the small $|\mu|^{2}$. The low energy physics is controlled by the $\Phi(r^{\mu})\sim\phi(r^{\mu})$ wave-function spreading, rather than the renormalization group of BHL . The hierarchy is protected against additive radiative corrections by the bilocality, i.e., there is no “additive quadratic divergence,” but only additive and enhancing renormalizations of the bilocal binding interaction chris2 . We used the source/Legendre-transform methods of Jackiw et.al., Jackiw , to derive the effective semiclassical theory used here, which leads to critical amplification of the potential coupling $\overline{g}_{0}^{2}$ chris2 . The fine-tuning is at the few $\%$ level, and, indeed, this may be the first and only mnimally-fine-tuned theory of the BEH boson that is consistent with experiment and testable in the not-too distant future.
Appendix A Brief Summary of Natural Top Condensation
@@
Our new formalism, “Natural Top Condensation,” chris1 ; chris2 , is Lorentz invariant and postulates an attractive “topcolor” interaction Topcolor of strength $g_{0}^{2}$ at a high scale $M_{0}$. This yields a bound state given by a Schrödinger-Klein-Gordon (SKG) equation satisfied by an internal wave-function, $\phi(r)$. This has eigenvalue $\mu^{2}$, which is the Lagrangian mass of the BEH boson and we have a Yukawa interaction between the bound state and unbound fermions. For supercritical coupling, $g_{0}>g_{c}$ we find $\mu^{2}<0$, and which implies spontaneous symmetry breaking. For small $|\mu|<M_{0}$, near critical coupling, we have significant wave-function spreading and “dilution” of $\phi(0)\sim\sqrt{|\mu|/M_{0}}$. The resulting top quark Yukawa, $g_{Y}\propto\phi(0)$, and quartic couplings, $\lambda\propto|\phi(0)|^{4}$, are subject to power law suppression, rather than the relatively slow renormalization group (RG) evolution in the old Nambu–Jona-Lasinio (NJL NJL ) based top condensation model. The dilution effect significantly reduces the hierarchy and, remarkably, the standard model (SM) quartic coupling, $\lambda\approx 0.25$, becomes concordant with experiment. The fine tuning of the model is also vastly reduced by dilution to $\sim\phi(0)^{2}\sim|\mu|/M_{0}\sim$ few $\%$. Our central prediction is the existence of a binding interaction due to a color octet of massive gluon-like objects, called “colorons,” Topcolor NSD colorons , with mass $M_{0}\sim 6$ TeV. The colorons may be accessible to the LHC. Moreover, loop effects enhance the binding in the $0^{+}$ channel, and the requisite $\mu^{2}<0$ can occur for significantly weaker coloron coupling chris2 . The top condensation model of a composite BEH boson therefore becomes a compelling theory. Inputting the induced Yukawa coupling $g_{Y}\approx 1$ we obtain the resulting prediction $M_{0}\sim 6$ TeV. This construction was confirmed by applying the formal source/Legendre-transformation methods of Jackiw, et. al, chris2 Jackiw .
@@
With the pullback, $\phi(r)$ is a dimensionless complex scalar and has no gauge charges. The Wilson line pullback is essentially a low energy approximation for the electroweak interactions, but makes the effects of symmetry breaking transparent.
In the barycentric coordinates we have:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(Z|D_{H}H(X)|^{2}|\phi(r)|^{2}+Z|H(X)|^{2}|\partial_{r}\phi(r)|^{2}+2{g_{0}^{2}N_{c}}D_{F}(2r)|H^{\dagger}H||\phi(r))|^{2}\biggr)\!\!+S_{Y}+S_{\lambda}+S^{\prime}$ (75)
where the $H(X)$ kinetic term is canonical with the Lorentz invariant normalization of $\phi(r)$ is:
$1=M_{0}^{4}Z\int\!\!d^{4}r|\phi(r^{\mu})|^{2}$ (76)
and $Z$ as defined independently below. The Yukawa interaction is also generated at tree level,
$\!\!\!\!\!\!\!\!\!\!S_{Y}=\widehat{g}_{Y}M_{0}^{2}\int\!\!d^{4}Xd^{4}r\;[\overline{\psi}_{iL}(X\!+\!r)\psi_{R}(X\!-\!r)]_{f}D_{F}(2r)\;H^{i}(X)\phi(\vec{r}){+h.c.}.$ (77)
@@
$g_{Y}\approx g_{0}^{2}\sqrt{2N_{c}/J}\;\phi(0)\qquad\qquad\lambda\approx(g_{Y}^{4}-g_{Y}^{2}\lambda)\frac{N_{c}}{4\pi^{2}}\ln\left(\frac{M_{0}}{\mu}\right),$ (79)
@@
$S=\!\!\int\!\!d^{4}X\;\biggl(|D_{H}H(X)|^{2}+|H(X)|^{2}S_{\phi}+\ g_{Y}H^{i\dagger}(X)[\overline{\psi}_{R}(X)\psi_{iL}(X)+h.c.]_{f}-\frac{1}{2}\lambda(H^{\dagger}H)^{2}\biggr)+S^{\prime},$ (80)
@@
$\!\!\!\!\!\!\!\!\!\!S^{\prime}$ $=$ $\!\int\!d^{4}x\biggl([\overline{\psi}_{L}(x)iD\!\!\!\!/\!\,\,\psi_{L}(x)]_{f}+[\overline{\psi}_{R}(x)\;iD\!\!\!\!/\!\,\,\psi_{R}(x)]_{f}\biggr)+g_{0}^{2}\!\int\!d^{4}xd^{4}y\;[\overline{\psi}^{i}_{L}(x)\psi_{R}(y)]_{f}D_{F}(x-y)[\overline{\psi}_{R}(y)\psi_{iL}(x)]_{f}$ (82)
In $S^{\prime}$ we have free unbound fermions, $\psi_{Lif}\sim(t,b)_{L}$ and $\psi_{Rf}\sim t_{R}$ (the minimal model omits $b_{r}$ but this can be readily incorporated as in Topcolor ). Note that the internal field action $S_{\phi}$ is nested within the action for a conventional pointlike BEH boson, $H(X)$. $D_{F}(2r^{\mu})$ is the Feynman propagator function for the massive colorons:
@@
$\lambda\approx(g_{Y}^{4}-g_{Y}^{2}\lambda-[\lambda^{2}])\frac{N_{c}}{4\pi^{2}}\ln\left(\frac{M_{0}}{\mu}\right)\approx 0.23\;\;\;\makebox{(cf., $0.25$ experiment.)};$ (84)
@@
Appendix B Orthogonality
$1=ZM_{0}^{4}\!\!\int\!d^{4}r\;\phi_{\omega_{i}}^{\dagger}(r)\phi_{\omega_{i}}(r)=1=M_{0}^{3}\!\!\int\!(ZM_{0})\omega_{\mu i}dr^{\mu}\;d^{3}r_{\perp}\;\phi_{\omega_{i}}^{\dagger}(r)\phi_{\omega_{i}}(r)$ (85)
$1=M_{0}^{3}\!\!\int\!d^{3}r|\phi_{\omega}(\vec{r})|^{2}\qquad\text{where,}\qquad 1=ZM_{0}\int dr^{0}\equiv ZM_{0}T$ (86)
$1=M_{0}^{3}\!\!\int\!4\pi r^{2}dr\frac{N^{2}e^{-2|\mu|r}}{r^{2}}\sim 2\pi\frac{N^{2}M_{0}^{3}}{\mu}\qquad\text{hence,}\!\!\qquad\!\!N^{2}=\mu/2\pi M_{0}^{3}$ (87)
$\int d^{4}r\;\phi^{\dagger}_{\omega}(r)\phi_{\omega^{\prime}}(r)=0\qquad\omega_{\mu}\neq\omega^{\prime}_{\mu}$ (88)
$\phi_{\omega^{\prime}}(r^{\mu})=\phi(\omega^{\prime\mu}(\omega^{\prime}_{\nu}r^{\nu}/\omega^{\prime 2})-r^{\mu})=\phi(r^{0}\sinh^{2}\theta,r_{x}\cosh^{2}\theta,r_{y},r_{z})$ (89)
$\approx\frac{\pi Z{N^{2}}M^{4}_{0}}{\theta\mu^{2}}\ln(M_{0}T)=\frac{1}{2\mu\theta T}\ln(M_{0}T)$ (90)
In the $T\rightarrow\infty$ limit this approaches zero.
@@ Acknowledgments
I thank Bill Bardeen, Bogdan Dobrescu, Tao Han and other participants of Photon-Lepton 2025, for discussions and comments.

### v3

@@ V Summary
We have proposed a nontrivial vacuum for the natural top condensation theory. The vacuum is manifestly Lorentz invariant, composed of a collective Lorentz invariant sum over internal wave-functions, $\Phi=(1/\sqrt{N})\sum_{\omega}\phi_{\omega}(r^{\mu})$. Though each internal wave-function, $\phi_{\omega}$, is independent of relative time, $\tau$, they each have dependence upon the arrow of relative time, $\omega^{\mu}$, and are solutions to the Schrödinger-Klein-Gordon equation with eigenvalue $|\mu^{2}|$. The sum over $\phi_{\omega}$ makes the collective field, $\Phi(r^{\mu})$, Lorenz invariant and completely independent of $\omega^{\mu}$.
@@
$H^{\prime}(X)\Phi(r^{\mu})=H(X^{\mu})\widetilde{\Phi}(r)\rightarrow\exp(i\pi^{a}(X)\tau^{a}/2v_{weak})\biggl(\begin{array}[]{c}v_{weak}\widetilde{\Phi}(r^{\mu})+h(X,r^{\mu})\\ 0\end{array}\biggr),$ (72)
where $v_{weak}=|\mu|/\sqrt{2}$ and the observed BEH boson is $h(X)$ where:
$h(X,r^{\mu})=h(X)\widetilde{\Phi}(r^{\mu}).$ (73)
@@
Note that bilocality of the wave-function is an important naturalness constraint. One might be tempted to, e.g., “loop” the Yukawa interaction and argue for a problematic large correction to the BEH mass $\propto-\overline{g}_{Y}^{2}N_{c}M_{0}^{2}$. This would lead to the false conclusion that the effective theory is “unnatural.” The loop actually generates an enhancing correction to the bilocal potential, i.e., $\sim g_{Y}^{2}\delta^{4}(2r^{\mu})/M_{0}^{2}\sim g_{Y}^{2}D_{F}(2r^{\mu})$ for large $M_{0}$ chris2 , that leads to a “critical amplification” of the effective 4-fermion coupling, $\overline{g}^{2}_{0}$, where: $\overline{g}_{0}^{2}=g_{0}^{2}\left(1-{g_{0}^{2}N_{c}}/{8\pi^{2}}\right)^{-1}.$ Hence, while $\overline{g}_{0}^{2}$ is supercritical, the underlying topcolor coupling, $g_{0}^{2}$, is smaller and subcritical.
@@
In conclusion, we emphasize this theory is natural and manifestly Lorentz invariant. The approximate scale symmetry near critical coupling provides the custodial symmetry of the small $|\mu|^{2}$. The low energy physics is controlled by the $\Phi(r^{\mu})\sim\phi(r^{\mu})$ wave-function spreading, rather than the renormalization group of BHL . The hierarchy is protected against additive radiative corrections by the bilocality, i.e., there is no “additive quadratic divergence,” but only additive and enhancing renormalizations of the bilocal binding interaction chris2 . We used the source/Legendre-transform methods of Jackiw et.al., Jackiw , to derive the effective semiclassical theory used here, which leads to critical amplification of the potential coupling $\overline{g}_{0}^{2}$ chris2 . The fine-tuning is at the few $\%$ level, and, indeed, this may be the first and only minimally-fine-tuned theory of the BEH boson that is consistent with experiment and testable in the not-too distant future.
Appendix A Orthogonality of $\phi_{\omega}(r)$
$1=ZM_{0}^{4}\!\!\int\!d^{4}r\;\phi_{\omega_{i}}^{\dagger}(r)\phi_{\omega_{i}}(r)=1=M_{0}^{3}\!\!\int\!(ZM_{0})\omega_{\mu i}dr^{\mu}\;d^{3}r_{\perp}\;\phi_{\omega_{i}}^{\dagger}(r)\phi_{\omega_{i}}(r).$ (74)
$1=M_{0}^{3}\!\!\int\!d^{3}r|\phi_{\omega}(\vec{r})|^{2},\qquad\text{where,}\qquad 1=ZM_{0}\int dr^{0}\equiv ZM_{0}T.$ (75)
$1=M_{0}^{3}\!\!\int\!4\pi r^{2}dr\frac{N^{2}e^{-2|\mu|r}}{r^{2}}\sim 2\pi\frac{N^{2}M_{0}^{3}}{\mu},\qquad\text{hence,}\!\!\qquad\!\!N^{2}=\mu/2\pi M_{0}^{3}.$ (76)
$\int d^{4}r\;\phi^{\dagger}_{\omega}(r)\phi_{\omega^{\prime}}(r)=0,\qquad\omega_{\mu}\neq\omega^{\prime}_{\mu}.$ (77)
$\phi_{\omega^{\prime}}(r^{\mu})=\phi(\omega^{\prime\mu}\omega^{\prime}_{\nu}r^{\nu}-r^{\mu})=\phi(r^{0}\sinh^{2}\theta,r_{x}\cosh^{2}\theta,r_{y},r_{z}).$ (78)
$\approx\frac{\pi Z{N^{2}}M^{4}_{0}}{\theta\mu^{2}}\ln(M_{0}T)=\frac{1}{2\mu\theta T}\ln(M_{0}T).$ (79)
In the $T\rightarrow\infty$ limit this approaches zero.
Appendix B Brief Summary of Natural Top Condensation
@@
Our formalism, “Natural Top Condensation,” chris1 ; chris2 , is Lorentz invariant and postulates an attractive “topcolor” interaction Topcolor of strength $g_{0}^{2}$ at a high scale $M_{0}$. The bound states are correlated pairs $\overline{\psi}(y)_{L}\psi_{R}(x)\rightarrow\Phi(x,y)\sim\Phi(X,r)$, and we follow Yukawa in writing invariant kinetic terms for the pairs, consistent with their single particle kinetic terms Yukawa . This yields a bound state given by a Schrödinger-Klein-Gordon (SKG) equation satisfied by an internal wave-function, $\phi(r)$. This has eigenvalue $\mu^{2}$, which is the Lagrangian mass of the BEH boson and we have a Yukawa interaction between the bound state and unbound fermions. For supercritical coupling, $g_{0}>g_{c}$ we find $\mu^{2}<0$, and which implies spontaneous symmetry breaking. For small $|\mu|<M_{0}$, near critical coupling, we have significant wave-function spreading and “dilution” of $\phi(0)\sim\sqrt{|\mu|/M_{0}}$. The resulting top quark Yukawa, $g_{Y}\propto\phi(0)$, and quartic couplings, $\lambda\propto|\phi(0)|^{4}$, are subject to power law suppression, rather than the relatively slow renormalization group (RG) evolution in the old Nambu–Jona-Lasinio (NJL NJL ) based top condensation model. The dilution effect significantly reduces the hierarchy and, remarkably, the standard model (SM) quartic coupling, $\lambda\approx 0.25$, becomes concordant with experiment. The fine tuning of the model is also vastly reduced by dilution to $\sim\phi(0)^{2}\sim|\mu|/M_{0}\sim$ few $\%$. Our central prediction is the existence of a binding interaction due to a color octet of massive gluon-like objects, called “colorons,” Topcolor NSD colorons , with mass $M_{0}\sim 6$ TeV. The colorons may be accessible to the LHC. Moreover, loop effects enhance the binding in the $0^{+}$ channel, and the requisite $\mu^{2}<0$ can occur for significantly weaker coloron coupling chris2 . The top condensation model of a composite BEH boson therefore becomes a compelling theory. Inputting the induced Yukawa coupling $g_{Y}\approx 1$ we obtain the resulting prediction $M_{0}\sim 6$ TeV. This construction was confirmed by applying the formal source/Legendre-transformation methods of Jackiw, et. al, chris2 Jackiw .
@@
With the pullback, $\phi(r)$ is a dimensionless complex scalar and has no gauge charges. The Wilson line pullback is essentially a low energy approximation for the electroweak interactions, but makes the effects of symmetry breaking transparent. - In the barycentric coordinates we have:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(Z|D_{H}H(X^{\mu})|^{2}|\phi(r)|^{2}+Z|H(X^{\mu})|^{2}|\partial_{r}\phi(r)|^{2}+2{g_{0}^{2}N_{c}}D_{F}(2r)|H^{\dagger}H||\phi(r))|^{2}\biggr)\!\!+S_{Y}+S_{\lambda}+$... (81)
where the $H(X^{\mu})$ kinetic term is canonical with the Lorentz invariant normalization of $\phi(r)$ is:
$1=M_{0}^{4}Z\int\!\!d^{4}r|\phi(r^{\mu})|^{2}\rightarrow M_{0}^{3}\int\!\!d^{3}r|\phi(r^{\mu})|^{2},$ (82)
where $Z$ is as defined upon removal of relative time as in (12). The Yukawa interaction is also generated at tree level,
$\!\!\!\!\!\!\!\!\!\!S_{Y}=\widehat{g}_{Y}M_{0}^{2}\int\!\!d^{4}Xd^{4}r\;[\overline{\psi}_{iL}(X\!+\!r)\psi_{R}(X\!-\!r)]_{f}D_{F}(2r)\;H^{i}(X)\phi(\vec{r}){+h.c.},$ (83)
@@
$g_{Y}\approx g_{0}^{2}\sqrt{2N_{c}/J}\;\phi(0),\qquad\qquad\lambda\approx(g_{Y}^{4}-g_{Y}^{2}\lambda)\frac{N_{c}}{4\pi^{2}}\ln\left(\frac{M_{0}}{\mu}\right).$ (85)
@@
$S=\!\!\int\!\!d^{4}X\;\biggl(|D_{H}H(X^{\mu})|^{2}+|H(X^{\mu})|^{2}S_{\phi}+\ g_{Y}H^{i\dagger}(X)[\overline{\psi}_{R}(X)\psi_{iL}(X)+h.c.]_{f}-\frac{1}{2}\lambda(H^{\dagger}H)^{2}\biggr)+S^{\prime},$ (86)
@@
$\!\!\!\!\!\!\!\!\!\!S^{\prime}$ $=$ $\!\int\!d^{4}x\biggl([\overline{\psi}_{L}(x)iD\!\!\!\!/\!\,\,\psi_{L}(x)]_{f}+[\overline{\psi}_{R}(x)\;iD\!\!\!\!/\!\,\,\psi_{R}(x)]_{f}\biggr)+g_{0}^{2}\!\int\!d^{4}xd^{4}y\;[\overline{\psi}^{i}_{L}(x)\psi_{R}(y)]_{f}D_{F}(x-y)[\overline{\psi}_{R}(y)\psi_{iL}(x)]_{f}.$ (88)
In $S^{\prime}$ we have free unbound fermions, $\psi_{Lif}\sim(t,b)_{L}$ and $\psi_{Rf}\sim t_{R}$ (the minimal model omits $b_{r}$ but this can be readily incorporated as in Topcolor ). Note that the internal field action $S_{\phi}$ is nested within the action for a conventional pointlike BEH boson, $H(X^{\mu})$. $D_{F}(2r^{\mu})$ is the Feynman propagator function for the massive colorons:
@@
$\lambda\approx(g_{Y}^{4}-g_{Y}^{2}\lambda-[\lambda^{2}])\frac{N_{c}}{4\pi^{2}}\ln\left(\frac{M_{0}}{\mu}\right)\approx 0.23\;\;\;\makebox{(cf., $0.25$ experiment.)},$ (90)
@@ Acknowledgments
I thank Bill Bardeen, Bogdan Dobrescu, Tao Han and other participants of Photon-Lepton 2025, for discussions and comments. Sadly, Bill Bardeen passed away during the completion of this work. His awesome intellect, depth of understanding of physics, and close friendship, are sorely missed.